40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3437     Entry time: Wed Aug 18 19:19:38 2010
Author: Jenne 
Type: Update 
Category: SUS 
Subject: Final 2 TTs suspended! 

[Jenne, Yoichi]

The final 2 Tip Tilts (#1 and #5) have been suspended.  We have designated #5 the spare.  It looks like there might be a teensy bit of dust on the AR surface of the optic in #5, right near the edge of the coating.  Not a critical issue if this one is the spare, although we should see if we can blow it off with the Nitrogen.  Both #1 and #5's optics were suspended using the thicker wire, 0.0036" diameter.  This leaves 4/5 TTs with this thick wire, and 1 of the 5 has the thin wire.

To do still: Balance both #1 and #5, and then measure the modes of each.  Then we'll be ready to install them into the chambers, and we'll reserve #5 for shake table TFs for some later date.

 

ELOG V3.1.3-